9. The area of a rectangle is 45 square cm. If the length is 4 cm greater than the width, what are the dimension of the rectangle?


Found 2 solutions by Molly27, checkley77:
Answer by Molly27(1)

Length is 9 width is 5

Answer by checkley77(12844)

L=W+4
AREA=LW
45=(W+4)W
45=W^2+4W
W^2+4W-45=0
(W+9)(W-5)=0
W-5=0
W=5 ANS FOR THE WIDTH.
L=5+4=9 FOR THE LENGTH.
PROOF
5*9=45